OpenStudy (anonymous):

FAHRENHEIT TO CELSIUS. Temperatures in degrees Celsius (C) and degrees Fahrenheit (F) are related by the formula: C = (5F-160) / 9 What is the temperature in degrees Celsius if it is 77 degrees F? A. 8° C B. 25° C C. 117° C D. answer not given

OpenStudy (anonymous):

first list down what u got here C = (5F-160) / 9 u also know F=77 now plug the value C=(5*77-160)/9=25 so the answer is B